OpenStudy (anonymous):

What is the equation of the line that passes through the points (0, 5) and (2, 3)?

OpenStudy (anonymous):

See, equation of line that passes through two points that are (x1, y1) and (x2, y2) is given by two point form: \[\color{green}{\large \frac{y - y_1}{y_2 - y_1} = \frac{x- x_1}{x_2 - x_1}}\]

OpenStudy (anonymous):

Here you have: \(x_1 = 0, x_2 = 2, y_1 = 5, y_2 = 3\) Just plug in these values in the above said equation..

OpenStudy (anonymous):

wha the answer

OpenStudy (anonymous):

I am afraid, I can't give you answer directly.. I have provided you all the necessary steps, just try it out, if you face any difficulty, just let me know, but first work it out on your own.. Just try, I am sure you can do it..

OpenStudy (anonymous):

i have difficulty

OpenStudy (anonymous):

And please, use this site for Learning purposes rather than asking direct answers.. Please invest some of your time to read the following link: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Yes, It is good if you have difficulty, just tell me where are you facing that difficulty??

OpenStudy (anonymous):

pluging it in

OpenStudy (anonymous):

Are you seeing that formula above I have written??

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

yesssss

OpenStudy (anonymous):

There in place of \(x_1) there just write 0..

OpenStudy (anonymous):

is the answer x+1?

OpenStudy (anonymous):

I mean just write there: \(x_1\) = 0 \(x_2\) = 2 \(y_1\) = 5 \(y_2\) = 3

OpenStudy (anonymous):

Why are you worrying for answer?? If you know how to solve, then answer is not far to achieve..

OpenStudy (anonymous):

Just try, you can do all wrong at the most, nothing greater than that..

OpenStudy (anonymous):

the answer is 2

OpenStudy (anonymous):

Okay. I accept you but I don't tell you what the answer is...